OpenStudy (anonymous):

You research the average cost of a tank of gasoline for a certain type of car for several recent years to look for trends. The table shows your data. What is the equation for a line of best fit? How much would you expect to pay for a tank of gas in the year 2019? Let x be the number of years after 1998.

OpenStudy (anonymous):

Year 1998 2000 2002 2004 2006 2008 Average cost for one tank ($) $26.40 $27.20 $29.70 $29.30 $32.20 $37.80

OpenStudy (anonymous):

A y = 1.006x + 26.40; $47.52 B y = 1.006x + 26.40; $58.92 C y = 26.40x + 1.006; $47.52 D y = 26.40x + 1.006; $58.92

OpenStudy (anonymous):

 Gradient : we need 2 coordinates ... for example : ( 29.30 , 2008 ) and ( 26.40 , 1998 ) Gradient , m : Y - y1 divide X -x1 can you carry on now ?